# RLSA-PFL: Robust Lightweight Secure Aggregation with Model Inconsistency Detection in Privacy-Preserving Federated Learning

### v2

@@ 4.4 Goals
Our Proposed Secure Aggregation Scheme Setup All the entities agree on the security parameter $\lambda$, a pseudo-random function $\mathrm{PRF}$, a hash function $H:\{0,1\}^{*}\rightarrow\{0,1\}^{l}$, a set of integer modulo $q$ $\mathbb{Z}_{q}$, and a threshold value $t$. Each user, intermediate server, and aggregator are assigned unique identifiers (denoted as $u_{i}$ for the $i^{th}$ user, $f_{i}$ for the $i^{th}$ intermediate server, and $\mathrm{Agg}$ for the aggregator). Each user, intermediate server, and aggregator are assigned a public and private key pair $(\mathsf{pub_{u_{i}}},\mathsf{priv_{u_{i}}}),(\mathsf{pub_{f_{i}}},\mathsf{priv_{f_{i}}})$, and $(\mathsf{pub_{Agg}},\mathsf{priv_{Agg}})$, respectively, for signature generation. All users know the participating intermediate servers and the aggregator set $\mathcal{N}$, where $|\mathcal{N}|=n$. All intermediate servers and the aggregator have the universal set of the users $\mathcal{M}$, where a random subset $\mathcal{U}$ ($m=|\mathcal{U}|$) of users participate in each round of training. All intermediate servers communicate with the aggregator over private, authenticated channels. All users have private, authenticated channels with both the intermediate servers and the aggregator. Masking Round Users: * Each user $u_{i}$ chooses random numbers $\{r_{j}\}_{\forall j\in\mathcal{N}}\in\mathbb{Z}_{q}$, and then generates the mask vectors $\{k_{j}=\mathsf{PRF}(r_{j})\}_{\forall j\in\mathcal{N}}$. * Each user $u_{i}$ generates a directed cycle for the participating entities in $\mathcal{N}$, as described in Section 5.1. * For each $j^{th}$ node in the directed cycle, each user $u_{i}$ masks the trained model $x_{t}$ as follows: $c^{i,j}_{t}=\frac{x_{t}}{n}+k_{j}-k_{j-1}\mod{q}$ and sends $<c^{i,j}_{t},\sigma^{i,j}_{t}>$ to the $j^{th}$ node, where $\sigma^{i,j}_{t}=\mathsf{SIG.Sign}(\mathsf{priv_{u_{i}}},c^{i,j}_{t})$. Intermediate Servers: * Each intermediate server, say $f_{j}$ chooses an empty list $\mathrm{F_{j}}$. * Each intermediate server $f_{j}$ receives the tuple $<c^{i,j}_{t},\sigma^{i,j}_{t}>$ from each participating user $u_{i}$. * The intermediate server computes $\mathsf{SIG.Ver}(\mathsf{pub_{u_{i}}},c^{i,j}_{t},\sigma^{i,j}_{t})$. If the signature is valid, it adds the user identity $u_{i}$ to the participating user list $\mathrm{F_{j}}$. * Each intermediate server $f_{j}$ sends the tuple $<\mathrm{F_{j}},\sigma_{\mathrm{F_{i}}}>$ to the aggregator if and only if $|\mathrm{F_{j}}|\geq t$, where $\sigma_{\mathrm{F_{j}}}=\mathsf{SIG.Sign}(\mathsf{priv_{f_{j}}},\mathrm{F_{j}})$. Otherwise, it aborts. Aggregator $\mathsf{Agg}$: * The aggregator chooses an empty list $\mathrm{A}$. * The aggregator receives the tuple $<c^{i,\mathsf{Agg}}_{t},\sigma^{i,\mathsf{Agg}}_{t}>$ from each user $u_{i}$. * It performs $\mathsf{SIG.Ver}(\mathsf{pub_{u_{i}}},c^{i,\mathsf{Agg}}_{t},\sigma^{i,\mathsf{Agg}}_{t})$. If the signature is valid, it adds the user identity $u_{i}$ to the participating user list $\mathrm{A}$. * If $|\mathrm{A}|\geq t$ and the aggregator receives all tuples <$<\{\mathrm{F_{j}},\sigma_{\mathrm{F_{j}}}\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}>$ from all the intermediate servers, it verifies the signatures $\{\mathsf{SIG.Ver}(\mathsf{pub_{f_{j}}},\mathrm{F_{j}},\sigma_{\mathrm{F_{j}}})\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}$; otherwise it aborts. * The aggregator computes the common active user list $\mathrm{I}=\mathrm{A}\cap\{\mathrm{F_{j}}\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}$. * If $\mathrm{|I|}\geq t$, the aggregator sends back the tuple $<\mathrm{I},\sigma_{\mathsf{Agg}}>$ to each intermediate server $f_{j}$, where $\sigma_{\mathsf{Agg}}=\mathsf{SIG.Sign(\mathsf{priv_{Agg}},I)}$. Partial Aggregation by the Intermediate Servers Each participating intermediate server, say $f_{j}$ receives the tuple $<\mathrm{I},\sigma_{\mathsf{Agg}}>$ from the aggregator. If $|\mathrm{I}|\geq t$, the intermediate server $f_{j}$ checks $\mathsf{SIG.Ver}(\mathsf{pub_{Agg}},\mathrm{I},\sigma_{\mathsf{Agg}})$. If the signature is valid, the intermediate server ${f_{j}}$ computes the partially aggregated masked model $\mathsf{PartialAgg_{f_{j}}}$ using the masked model received from the users in $\mathrm{I}$. $\mathsf{PartialAgg_{f_{j}}}=\sum_{\forall i\in\mathrm{I}}c^{i,j}_{t}$ The intermediate server $f_{j}$ sends the tuple $<\mathsf{PartialAgg_{f_{j}}},\sigma_{\mathsf{PartialAgg_{f_{j}}}}>$ to the aggregator, where $\sigma_{\mathsf{PartialAgg_{f_{j}}}}=\mathsf{SIG.Sign}(\mathsf{priv_{f_{j}}},\mathsf{PartialAgg_{f_{j}}})$. Final Aggregation and Unmasking by the Aggregator Once the aggregator receives all the tuples $\{\mathsf{PartialAgg_{f_{j}}},\sigma_{\mathsf{PartialAgg_{f_{j}}}}\}_{\forall f_{i}\in(\mathcal{N}\setminus\mathrm{Agg})}$, it checks $\{\mathsf{SIG.Ver}(\mathsf{pub_{f_{j}}},\mathsf{PartialAgg_{f_{j}}},\sigma_{\mathsf{PartialAgg_{f_{j}}}})\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}$. If the signatures are valid, the aggregator performs the final aggregation as follows: $\theta_{t}=\sum_{\forall i\in\mathrm{I}}c^{i,\mathrm{Agg}}_{t}+\sum_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}\mathsf{PartialAgg_{f_{j}}}$ Aggregator sends the global model parameter $\theta_{t}$ and the common active user list $\mathrm{I}$ to each participating user for the next round of training. Global Model Parameter Verification Aggregator: * Aggregator chooses a random number $s_{t}\in\mathbb{Z}_{q}$. * Aggregator computes $R=H(\theta_{t})+s_{t}$ and a message authentication code $S=\mathtt{MAC}_{s_{t}}(\theta_{t})$ Aggregator sends the tuple $V=<\mathrm{T}=<R,S>,\mathrm{A},\mathsf{SIG.Sign(\mathsf{priv_{Agg}},<T,\mathrm{I,A}>)}>$ to each intermediate server. Intermediate Servers: * Each intermediate server, say $f_{j}$ verifies $\mathsf{SIG.Sign(\mathsf{priv_{Agg}},<\mathrm{T},\mathrm{I,A}>)}$ once it received the tuple $<\mathrm{T}=<R,S>,\mathrm{A},\mathsf{SIG.Sign(\mathsf{priv_{Agg}},<T,\mathrm{I,A}>)}>$ from the aggregator. * If the verification is successful, the intermediate server $f_{j}$ forwards the tuple $<\mathrm{T}$, $\mathrm{I,A}>$ to the users in $\mathrm{I}$ along with the user list $\mathrm{F_{j}}$. User: * Each user $u_{i}$ receives the tuple $<\mathrm{T}=<R,S>,\mathrm{I,A}>$ and the user list $\{\mathrm{F_{j}}\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}$ from the intermediate servers. * The user first verifies if $\mathrm{I}=\mathrm{A}\cap\{\mathrm{F_{j}}\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}$ and $|\mathrm{I}|\geq t$. If verification fails, the user stops from participating in future rounds. Otherwise, the user goes to the following steps. * The user $u_{i}$ recovers the secret key $s_{t}=R-H(\theta_{t})$ (where the user already has $\theta$ from the aggregator). * If $S=\mathtt{MAC_{s_{i}}}(\theta_{t})$, the user $u_{i}$ compares the remaining $S$ values received from the other intermediate servers to verify their consistency. If the verification fails or any mismatches are found, the user $u_{i}$ detects a model inconsistency attack and ceases participation.
Figure 3: Detailed Description of Our Proposed Secure Aggregation Scheme
5 Our Proposed Scheme
@@
In this section, we provide a detailed explanation of our scheme. We first present the technical intuition behind the scheme, followed by a description of its main construction.
5.1 Technical Intuition
@@
Our protocol has three main entities: users, intermediate servers, and the aggregator, as described in Section 4.1. Users mask their locally trained models, intermediate servers partially aggregate these masked models, and the aggregator completes the final aggregation, eventually unmasking the global model. In this section, we explain how masking and unmasking work, along with how our scheme manages user dropouts and detects model inconsistency. Let $x_{u}$ be the private vectors of a user $u$, and $\mathcal{U}$ be the set of participating users in a round.
@@
The main goal of our scheme is to compute $\sum_{\forall u\in\mathcal{U}}x_{u}$ in such a way that the aggregator cannot see the individual private vectors $x_{u}$ of any user. Our scheme can tolerate up to $m-2$ compromised users, where $m=|\mathcal{U}|$. This means that the aggregator will only get at most the aggregated vectors of the two honest users, without learning their individual private vectors. Moreover, our scheme ensures that no user learns any useful information about another user’s private vector.
@@
In our system, there are $m$ participating users, $n$ intermediate servers along with the aggregator. We represent the set of intermediate servers and the aggregator as $\mathcal{N}$. Each user $u$ generates a random number $r_{i}$ in $\mathbb{Z}_{q}$ for each $i\in\mathcal{N}$. The user then uses a pseudo-random function $\mathsf{PRF}$ to generate a random vector $k_{i}=\mathsf{PRF}(r_{i})$ for each random number $r_{i}$ to mask its model. The user $u$ masks its model $x_{u}$ using the formula:
$c^{i}=$ $x_{u}+k_{i}\mod{q}.$ (6)
@@
The masking process described earlier does not involve an unmasking step, meaning the aggregator only receives the combined masked models. To introduce the unmasking step, our scheme uses a similar key negation approach as in [37], which is based on the aggregated homomorphic encryption method outlined in Equation 4 (see Section 3.1). The main goal is to use two random secret vectors to mask the models for each element in $\mathcal{N}$ so that these vectors cancel each other out during aggregation.
@@
[…]
$\mathtt{k_{i+2}}]\mod{q}.$ (7)
@@
$\frac{\left(\sum_{1\leq i\leq n}c_{i}\right)}{n}=$ $n\cdot x_{1}+n\cdot x_{2}+\cdots+n\cdot x_{u}+$
[…]
$=$ $\sum_{1\leq u\leq n}x_{u}.$ (8)
@@
Before describing our user dropout mechanism, we briefly explain the one used in the BBGLR protocol. In the BBGLR protocol and its variants, users agree on shared secret values with each other. This ensures that their random masks cancel out during the aggregation process. If a user drops out, the aggregator must still be able to remove that user’s masked contribution. To do this, the aggregator collects the secret shares from the remaining users to cancel out the dropped user’s contribution in the final aggregation. A similar method is also used in the Flamingo protocol [19], where the aggregator must contact a set of decryptors to handle the dropped users’ contributions.
In contrast, our scheme does not require users to agree on shared secrets or interact with each other. Each user’s masked model is independent of the others, unlike the BBGLR protocol. Our dropout mechanism is straightforward and requires only an additional communication round between the intermediate servers and the aggregator. After the masking phase, each user sends their masked models to both the intermediate servers and the aggregator. Consequently, each intermediate server $f_{i}$ and the aggregator generate a user list, $\mathrm{F_{i}}$ and $\mathrm{A}$ respectively, based on the users from whom they received masked models. Finally, the aggregator obtains the common active user list, $\mathrm{A}$ by collecting the user lists $\mathrm{F_{i}}$ from all intermediate servers. This list represents the active users for that round, from whom all entities received a masked model. All the intermediate servers and the aggregator use the user list $\mathsf{I}$ to aggregate the masked models. This process ensures that dropped users are identified, and their contributions are excluded from the aggregation without compromising the global model.
@@
In [20], it is explained how a malicious server (or the aggregator in our case) can carry out a model inconsistency attack. In this type of attack, as we explained earlier, the server manipulates the global model to provide different versions to specific users, potentially exposing sensitive information about them. In [20], one of the suggested ways to prevent issues is to make sure all users get the exact same global model from the aggregator. Our scheme uses this approach too.
In our scheme, the aggregator sends the global model along with a MAC to the intermediate servers. These servers then pass the MACs to all active users listed in $\mathsf{I}$. Since each user gets the same MACs from every intermediate server, they can check for mismatches. If a user receives any different MACs, they stop participating in the training process, as this indicates a model inconsistency attack.
5.2 Main Scheme
Our scheme involves $m$ users and $n$ intermediate servers, along with an aggregator. Each user $u_{i}\in\mathcal{U}$ holds a private vector $x_{t}$ of size $n$ for the round $t$, where each element of $x_{t}$ belongs to $\mathbb{Z}_{q}$ for some $q$. Both $m$ and $n$ are polynomially bounded. The $i^{th}$ user and $j^{th}$ intermediate server are denoted as $u_{i}$ and $f_{j}$, respectively. The security of our scheme relies on the initial security parameter $\lambda$, which defines the overall strength of the scheme. Our scheme also uses a PKI, assigning each entity a private and public key pair for signing messages in the malicious model. The detailed description of our proposed secure aggregation scheme is shown in Figure 3.

### v3

@@ 5.3. Assumption
Our Proposed Secure Aggregation Protocol Setup All the entities agree on the security parameter $\lambda$, a hash function $H:\{0,1\}^{*}\rightarrow\{0,1\}^{l}$, a set $\mathbb{Z}_{q}$ of integer modulo $q$, and a threshold value $t$. Each user, intermediate server, and aggregator are assigned unique identifiers (denoted as $u_{i}$ for the $i^{th}$ user, $f_{i}$ for the $i^{th}$ intermediate server, and $\mathrm{Agg}$ for the aggregator). Each user, intermediate server, and aggregator are assigned a public and private key pair $(\mathsf{pub_{u_{i}}},\mathsf{priv_{u_{i}}}),(\mathsf{pub_{f_{i}}},\mathsf{priv_{f_{i}}})$, and $(\mathsf{pub_{Agg}},\mathsf{priv_{Agg}})$, respectively, for signature generation. All users know the participating intermediate servers and the aggregator set $\mathcal{N}$, where $|\mathcal{N}|=n$. All intermediate servers and the aggregator have the universal set of the users $\mathcal{M}$, where a random subset $\mathcal{U}$ ($m=|\mathcal{U}|$) of users participate in each round of training. All intermediate servers communicate with the aggregator over private, authenticated channels. All users have private, authenticated channels with both the intermediate servers and the aggregator. Masking Round Users: * Each user $u_{i}$ chooses random masks (i.e., secret vectors) $\{k_{j}\}_{\forall j\in\mathcal{N}}\in\mathbb{Z}_{q}$. * Each user $u_{i}$ generates a directed cycle for the participating entities in $\mathcal{N}$, as described in Section 6.1. * For each $j^{th}$ node in the directed cycle, each user $u_{i}$ masks the trained model $x_{t}$ as follows: $c^{i,j}_{t}=\frac{x_{t}}{n}+k_{j}-k_{j-1}\mod{q}$ and sends $c^{i,j}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma^{i,j}_{t}}>$ to the $j^{th}$ node, where $\sigma^{i,j}_{t}=\mathsf{SIG.Sign}(\mathsf{priv_{u_{i}}},c^{i,j}_{t})$. Intermediate Servers: * Each intermediate server, say $f_{j}$ chooses an empty list $\mathrm{F_{j}}$. * Each intermediate server $f_{j}$ receives the tuple $c^{i,j}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma^{i,j}_{t}}>$ from each participating user $u_{i}$. * The intermediate server computes $\mathsf{SIG.Ver}(\mathsf{pub_{u_{i}}},c^{i,j}_{t},\sigma^{i,j}_{t})$. If the signature is valid, it adds the user identity $u_{i}$ to the participating user list $\mathrm{F_{j}}$. * Each intermediate server $f_{j}$ sends the tuple $\mathrm{F_{j}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma_{\mathrm{F_{i}}}}>$ to the aggregator if and only if $|\mathrm{F_{j}}|\geq t$, where $\sigma_{\mathrm{F_{j}}}=\mathsf{SIG.Sign}(\mathsf{priv_{f_{j}}},\mathrm{F_{j}})$. Otherwise, it aborts. Aggregator $\mathsf{Agg}$: * The aggregator chooses an empty list $\mathrm{A}$. * The aggregator receives the tuple $c^{i,\mathsf{Agg}}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma^{i,\mathsf{Agg}}_{t}}>$ from each user $u_{i}$. * It performs $\mathsf{SIG.Ver}(\mathsf{pub_{u_{i}}},c^{i,\mathsf{Agg}}_{t},\sigma^{i,\mathsf{Agg}}_{t})$. If the signature is valid, it adds the user identity $u_{i}$ to the participating user list $\mathrm{A}$. * If $|\mathrm{A}|\geq t$ and the aggregator receives all tuples $\{\mathrm{F_{j}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma_{\mathrm{F_{j}}}}\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}>$ from all the intermediate servers, it verifies the signatures $\{\mathsf{SIG.Ver}(\mathsf{pub_{f_{j}}},\mathrm{F_{j}},\sigma_{\mathrm{F_{j}}})\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}$; otherwise it aborts. * The aggregator computes the common active user list $\mathrm{I}=\mathrm{A}\cap\{\mathrm{F_{j}}\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}$. * If $\mathrm{|I|}\geq t$, the aggregator sends back the tuple $\mathrm{I},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma_{\mathsf{Agg}}}>$ to each intermediate server $f_{j}$, where $\sigma_{\mathsf{Agg}}=\mathsf{SIG.Sign(\mathsf{priv_{Agg}},I)}$. Partial Aggregation by the Intermediate Servers Each participating intermediate server, say $f_{j}$ receives the tuple $\mathrm{I},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma_{\mathsf{Agg}}}>$ from the aggregator. If $|\mathrm{I}|\geq t$, the intermediate server $f_{j}$ checks $\mathsf{SIG.Ver}(\mathsf{pub_{Agg}},\mathrm{I},\sigma_{\mathsf{Agg}})$. If the signature is valid, the intermediate server $f_{j}$} computes the partially aggregated masked model $\mathsf{PartialAgg_{f_{j}}}$ using the masked model received from the users in $\mathrm{I}$. $\mathsf{PartialAgg_{f_{j}}}=\sum_{\forall i\in\mathrm{I}}c^{i,j}_{t}$ The intermediate server $f_{j}$ sends the tuple $\mathsf{PartialAgg_{f_{j}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma_{\mathsf{PartialAgg_{f_{j}}}}}>$ to the aggregator, where $\sigma_{\mathsf{PartialAgg_{f_{j}}}}=\mathsf{SIG.Sign}(\mathsf{priv_{f_{j}}},\mathsf{PartialAgg_{f_{j}}})$. Final Aggregation and Unmasking by the Aggregator Once the aggregator receives all the tuples $\{\mathsf{PartialAgg_{f_{j}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma_{\mathsf{PartialAgg_{f_{j}}}}}\}_{\forall f_{i}\in(\mathcal{N}\setminus\mathrm{Agg})}$, it checks $\{\mathsf{SIG.Ver}(\mathsf{pub_{f_{j}}},\mathsf{PartialAgg_{f_{j}}},\sigma_{\mathsf{PartialAgg_{f_{j}}}})\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}$. If the signatures are valid, the aggregator performs the final aggregation as follows: $\theta_{t}=\sum_{\forall i\in\mathrm{I}}c^{i,\mathrm{Agg}}_{t}+\sum_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}\mathsf{PartialAgg_{f_{j}}}$ Aggregator sends the global model parameter $\theta_{t}$ and the common active user list $\mathrm{I}$ to each participating user for the next round of training. Global Model Parameter Verification Aggregator: * Aggregator chooses a random vector $s_{t}\in\mathbb{Z}_{q}$. * Aggregator computes $R=H(\theta_{t})+s_{t}$ and a message authentication code $S=\mathtt{MAC}_{s_{t}}(\theta_{t})$ * Aggregator sends the tuple $V=<\mathrm{T}=<R,S>,\mathrm{A},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathsf{SIG.Sign(\mathsf{priv_{Agg}},<T,\mathrm{I,A}>)}}>$ to each intermediate server. Intermediate Servers: * Each intermediate server, say $f_{j}$ verifies $\mathsf{SIG.Sign(\mathsf{priv_{Agg}},<\mathrm{T},\mathrm{I,A}>)}$ once it received the tuple $\mathrm{T}=<R,S>,\mathrm{A},\mathsf{SIG.Sign(\mathsf{priv_{Agg}},<T,\mathrm{I,A}>)}>$ from the aggregator. * If the verification is successful, the intermediate server $f_{j}$ forwards the tuple $\mathrm{T}$, $\mathrm{I,A}>$ to the users in $\mathrm{I}$ along with the user list $\mathrm{F_{j}}$. User: * Each user $u_{i}$ receives the tuple $\mathrm{T}=<R,S>,\mathrm{I,A},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathsf{SIG.Sign(\mathsf{priv_{Agg}},<T,\mathrm{I,A}>)}}>$ and the user list $\{\mathrm{F_{j}}\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}$ from the intermediate servers. * The user first verifies $\mathsf{SIG.Sign(\mathsf{priv_{Agg}},<T,\mathrm{I,A}>)}$. If verification is successful, it verifies if $\mathrm{I}=\mathrm{A}\cap\{\mathrm{F_{j}}\}_{\forall j\in(\mathcal{N}\setminus\mathrm{Agg})}$ and $|\mathrm{I}|\geq t$. If verification fails, the user stops from participating in future rounds. Otherwise, the user goes to the following steps. * The user $u_{i}$ recovers the secret key $s_{t}=R-H(\theta_{t})$ (where the user already has $\theta$ from the aggregator). * If $S=\mathtt{MAC_{s_{i}}}(\theta_{t})$, the user $u_{i}$ compares the remaining $S$ values received from the other intermediate servers to verify their consistency. If the verification fails or any mismatches are found, the user $u_{i}$ detects a model inconsistency attack and ceases participation.
Figure 3. Detailed Description of Our Proposed Secure Aggregation Protocol
6. Our Proposed Protocol
@@
In this section, we provide a detailed explanation of our proposed protocol. We first present the technical intuition behind the protocol, followed by a description of its main construction.
6.1. Technical Intuition
@@
Our protocol has three main entities: users, intermediate servers, and the aggregator, as described in Section 5.1. Users mask their locally trained models, intermediate servers partially aggregate these masked models, and the aggregator completes the final aggregation, eventually unmasking the global model. In this section, we explain how masking and unmasking work, along with how our protocol manages user dropouts and detects model inconsistency. Let $x_{u}$ be the private vectors of a user $u$, and $\mathcal{U}$ be the set of participating users in a round.
@@
The main goal of our protocol is to compute $\sum_{\forall u\in\mathcal{U}}x_{u}$ in such a way that the aggregator cannot see the individual private vectors $x_{u}$ of any user. Our protocol can tolerate up to m2m-2 compromised users, where $m=|\mathcal{U}|$. This means that the aggregator will only get at most the aggregated vectors of the two honest users, without learning their individual private vectors. Moreover, our protocol ensures that no user learns any useful information about another user’s private vector.
@@
In our system, there are $m$ participating users, $n$ intermediate servers along with the aggregator. We represent the set of intermediate servers and the aggregator as $\mathcal{N}$. Each user $u$ generates a random vector $k_{i}$ in $\mathbb{Z}_{q}$ for each $i\in\mathcal{N}$. The user $u$ masks its model $x_{u}$ using the formula:
(6) $c^{i}=$ $x_{u}+k_{i}\mod{q}.$
@@
The masking process described earlier does not involve an unmasking step, meaning the aggregator only receives the combined masked models. To introduce the unmasking step, our protocol uses a key negation approach, which is based on the aggregated homomorphic encryption method outlined in Equation 4 (see Section 4.1). The main goal is to use two random secret vectors to mask the models for each element in $\mathcal{N}$ so that these vectors cancel each other out during aggregation.
@@
$=$ $[x_{u}+x_{u+1}+\mathtt{k_{i}}-\cancel{\mathtt{k_{i+1}}}+\cancel{\mathtt{k_{i+1}}}$-
(7) $\mathtt{k_{i+2}}]\mod{q}.$
@@
$\frac{\left(\sum_{1\leq i\leq n}c_{i}\right)}{n}=$ $n\cdot x_{1}+n\cdot x_{2}+\cdots+n\cdot x_{u}+(\cancel{k_{1}}-\cancel{k_{2}})+(\cancel{k_{2}}-\cancel{k_{1}}+$
(8) $\cdots+(\cancel{k_{n-1}}-\cancel{k_{n}})+(\cancel{k_{n}}-\cancel{k_{1}}))=\sum_{1\leq u\leq n}x_{u}.$
@@
In the BBGLR protocol and its variants, users agree on shared secret values with each other. This ensures that their random masks cancel out during the aggregation process. If a user drops out, the aggregator must still be able to remove that user’s masked contribution. To do this, the aggregator collects the secret shares from the remaining users to cancel out the dropped user’s contribution in the final aggregation. A similar method is also used in the Flamingo protocol (Ma et al., 2023), where the aggregator must contact a set of decryptors to handle the dropped users’ contributions.
In contrast, our protocol does not require users to agree on shared secrets or interact with each other. Each user’s masked model is independent of the others, unlike the BBGLR protocol. Our dropout mechanism is straightforward and requires only an additional communication round between the intermediate servers and the aggregator. After the masking phase, each user sends their masked models to both the intermediate servers and the aggregator. Consequently, each intermediate server $f_{i}$ and the aggregator generate a user list, $\mathrm{F_{i}}$ and $\mathrm{A}$ respectively, based on the users from whom they received masked models. Finally, the aggregator obtains the common active user list, $\mathrm{I}$ by collecting the user lists $\mathrm{F_{i}}$ from all intermediate servers. This list represents the active users for that round, from whom all entities received a masked model. All the intermediate servers and the aggregator use the user list $\mathsf{I}$ to aggregate the masked models. This process ensures that dropped users are identified, and their contributions are excluded from the aggregation without compromising the global model.
@@
In (Pasquini et al., 2022), it is explained how a malicious server (or the aggregator in our case) can carry out a model inconsistency attack. In this type of attack, as we explained earlier, the server manipulates the global model to provide different versions to specific users, potentially exposing sensitive information about them. In (Pasquini et al., 2022), one of the suggested ways to prevent issues is to make sure all users get the exact same global model from the aggregator. Our protocol aims to use this technique.
In our protocol, the aggregator sends the global model along with a MAC to the intermediate servers. These servers forward the MACs to all active users in $\mathsf{I}$. Since each user receives the same MACs from every intermediate server, they can easily detect inconsistencies by checking for mismatches. If a user observes any discrepancy, they stop participating in the training process, as this signals a model inconsistency attack [^6] Our malicious model uses digital signatures to detect any tampering with MACs or other messages, preventing false accusations between participants..
Forward and Backward secrecy
One of the main limitations of e-SeaFL and Flamingo is their lack of FS and BS, due to reliance on a long-term secret seed as discussed in Appendix A. This seed is used to generate per-element masking values for each FL round via PRFs. If the seed is compromised, all associated past and future masking values can be recomputed. Our protocol avoids this problem by using fresh, random masking vectors from $\mathbb{Z}_{q}$ in every training round. These masks are never reused and are not derived from any long-term secret. As a result, even if an adversary learns the masks from one round, it does not affect the security of previous or future rounds. Each round is thus cryptographically independent, providing strong FS and BS guarantees.
Dynamic User Settings
As described earlier, our protocol does not require users to pre-establish shared secret seeds to participate in a training round, unlike existing protocols such as e-SeaFL and Flamingo. Instead, users can independently select fresh, random masking vectors from $\mathbb{Z}_{q}$ in each training round. These vectors are eventually canceled out during aggregation, thanks to our key negation technique and the intermediate server layer. There is no requirement to register users during the initial setup phase for key exchange. Under our security model, any permitted new user can join and leave the FL training at any round without compromising correctness or security.
6.2. Main Protocol
Our protocol involves $m$ users and $n$ intermediate servers, along with an aggregator. Each user $u_{i}\in\mathcal{U}$ holds a private vector $x_{t}$ of size $n$ for the round $t$, where each element of $x_{t}$ belongs to $\mathbb{Z}_{q}$ for some $q$. Both $m$ and $n$ are polynomially bounded. The $i^{th}$ user and $j^{th}$ intermediate server are denoted as $u_{i}$ and $f_{j}$, respectively. The security of our protocol relies on the initial security parameter $\lambda$, which defines the overall strength of the protocol. Our protocol also uses a PKI, assigning each entity a private and public key pair for signing messages in the malicious model. The detailed description of our proposed secure aggregation protocol is shown in Figure 3.